OpenStudy (cenaida):

If the equation of a quadratic is y = 6x2 -4x - 1, then the equation of the axis of symmetry is: A. B. C. 3 D. -3

OpenStudy (cenaida):

A. x=-1/3 B. x=1/3

OpenStudy (cenaida):

@SACAPUNTAS

OpenStudy (anonymous):

Do you know how to find the vertex?

OpenStudy (cenaida):

no

OpenStudy (anonymous):

Ok, the x coordinate of the vertex is \(\frac{-b}{2a}\) You can plug that back in if you wanted to find the y value but in this case that's all we need because: 

OpenStudy (cenaida):

so is it in between A and B?

OpenStudy (anonymous):

\[\frac{-b}{2a} = \frac{-(-4)}{2(6)} = \frac{4}{12} = \frac{1}{3}\]

OpenStudy (cenaida):

its B! thnks so much